12. Consider a spacecraft in an elliptical orbit around the Earth. At the low point, or perigee, of its orbit, it is 250 km aanswer Is b, please explain how to get to this
Answer #1

the velocity of perigee can be found as

v = sqrt ( 2 * GM * Ra / Rp * (Ra + Rp)

where M is mass of earth, Rp is perigee distance, Ra is apogee distance

remember, we are given distance from surface of earth so we have to account for radius.

Rp = 250000 + 6375000 = 6625000 m

Ra = 3500 + 6375 = 9875000 m

so,

v = sqrt ( 2 * 3.986e14 * 6625000 / 9875000 * ( 9875000 + 6625000 )

v = 5693.3 m/s

so,

the closest option is option (b)

so,

option (b) is correct.